OpenStudy (anonymous):

Simplify [ ( ½)^2 – ( ¼ )^3 ] x 2^3

OpenStudy (anonymous):

(1/2)^2 is the same as \[1^{2} \div 2^{2}= 1/4\] because you are multiplying the squared number by the top and bottom of the fraction. (1/4)^3 is the same as 1^3 divided by 4^3= 1/64 Then, use the order of operations.